if a 15.0 celcius tempature increase causes a 10.0 percent increase in the volume of a 832 ml sample of He(g) while the gas pressure is held constant what was the oridinal temperature?
1
Expert's answer
2019-10-11T09:36:55-0400

Use Charle's law



"\frac{V_1}{T_1}=\frac{V_2}{T_2}"

"T_2 = 15+T_1"

"V_1 = 832 cm^3"

"V_2 = 832 + 0.1\times832=915.2 cm^3"



"\frac{832}{T_1}=\frac{915.2}{T_1+15}"

"T_1 (K) = 150 K"

"T_2 (^\circ C) = 150 -273 = -123 =-123^\circ C"